(Y^2+yz+z^2)dx+(z^2+zx+x^2)dy+(x^2+xy+y^2)dz


1
Expert's answer
2021-01-04T17:22:58-0500

(y2+yz+z2)dx+(z2+zx+x2)dy+(x2+xy+y2)dz = (y2+z2)dx+(z2+x2)dy+(x2+y2)dz +d(xyz) = (x2+y2+z2)(dx+dy+dz) - (x2dx+y2dy+z2dz) - d(xyz) = (x2+y2+z2)d(x+y+z) - 1/3 d(x3+y3+z3 -3xyz)

As this form is symmetric with respect to the variables x, y and z, we should use the basic symmetric functions s=x+y+z, u=xy+yz+zx, v = xyz as new variables. We have then

x2+y2+z2=(x+y+z)2 - 2(xy+yz+zx) = s2-2u

x3+y3+z3 -3xyz = (x+y+z)3 - 3(x2y + xy2 + yz2 + y2z + z2x + zx2) - 9xyz = (x+y+z)3 - 3(x+y+z)(xy+yz+zx) = s3-3su

(x2+y2+z2)d(x+y+z) - 1/3 d(x3+y3+z3 -3xyz) = (s2-2u)ds - 1/3 d(s3-3su) = (s2-2u)ds - (s2-u)ds +s du = s du - u ds = s2 d(u/s) = -u2 d(s/u)

So, with the integrating factor s-2 = (x+y+z)-2 the original differential form becomes integrable, and general solution of the differential equation, determined by this differential form, in the domain U1 = {x+y+z not equal to 0} has the form f(u/s) = f((xy+yz+zx)/(x+y+z)), where f() is an arbitrary differentiable function.

In the domain U2 = {xy+yz+zx not equal to 0} general solution of the differential equation has the form g(s/u) = g((x+y+z)/(xy+yz+zx)), where g() is an arbitrary differentiable function. The integrating factor in this domain is u-2 = (xy+yz+zx)-2.